40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Dec 3 17:04:12 2012, Ayaka, Update, PSL, PMC calibration for MC_F calibration DSC_4955.JPGsweep_PZTdriver.pngPMC_PZT_FSR.pdftrans_fit.pngPMC-PZTcal_121203.zip
    Reply  Tue Dec 4 01:43:37 2012, rana, Update, PSL, PMC calibration for MC_F calibration 
    Reply  Tue Dec 4 11:30:36 2012, Zach, Update, PSL, PMC calibration for MC_F calibration 
       Reply  Thu Dec 13 17:24:53 2012, Ayaka, Update, PSL, PMC calibration for MC_F calibration 6x
Message ID: 7823     Entry time: Thu Dec 13 17:24:53 2012     In reply to: 7782
Author: Ayaka 
Type: Update 
Category: PSL 
Subject: PMC calibration for MC_F calibration 

 I calibrated MC_F signal into Hz/rtHz unit using the transfer function from MC_F to PMC feedback signal.

Here is the diagram:
  MC-PMCservo.jpg

n_mcf is MC_F signal we can get at dtt. I measured n_pmc/n'_mcf using SR.

TF_mcf-pmc.png

Other information I used:

G_out = 2.49/123.49 (see the document D980352-E01-C)

Fout has 1 pole at 10 Hz (see the document D980352-E01-C)

A_pzt = 371e+6/63 [Hz/V] (see elog)

F_wt has 1 pole at 100 Hz and 1 zero at 10 Hz.

Then, calibration transfer function of H is fitted as 1e+9/f [Hz/V]:

H.png

Then, the calibrated spectrum of MC_F is below:

 MCF_calib.pdf

This calibration have about 20 % error.
Compared to the spectrum in Jenne's paper (elog), above 20 Hz it seems to be laser frequency noise. But now we have extra unknown noise below 10 Hz.

Note: calibration value of laser's PZT is ~ 1MHz/V. This is reasonable compared to the data sheet of the laser. (This is calculated by combining result of H and transfer function of the circuit box1 and FSS.)

 laser.png

Attachment 6: calib.zip  31 kB
ELOG V3.1.3-